# Identifying the post-pandemic determinants of low performing students in Latin America through interpretable Machine Learning SHAP Values—Insights from PISA 2022

### v1

@@ 5.4 Countries’ results
To evaluate the possibility of heterogeneity of estimates of Sections 5.2 and 5.3, in this section I by examine SHAP values at the country level. Specifically, I run a similar analysis per country for the level 0 and level 1 sample8 yielding local explanations of SHAP values (students at extremes of the SHAP values distribution). But, first, I present the global contributions—average SHAP (absolute) values for all covariates.9
As a summary, I plot the frequency of the highest SHAP values contributions on predictions across the 10 countries. In Figure 6 I select the top 10 ranked covariates in terms of their SHAP average (absolute) values for each country and then obtain the number of times a given covariate appears in the top 10 list. Figure 6 shows that primary repetition appears in the top ranked list for all 10 countries; while family SES and school climate (in 9 countries) and digital devices (in 8 countries). This is followed by lack of school safety and homework intensity (present in 6 countries), whereas the covariates paid work, mother education and number of books appears in the top ten ranked covariates for 5 countries.
@@
As regards to local explanations (estimates shown in Tables 6 and 7) —where I run a full model with all the explanatory variables—, the main takeaways concerning leading drivers for students more likely to reach either level 0 or level 1 achievement per country are outlined below.10 For the first set of five countries (Table 6):
  • Argentina. (a) The student with the largest probability of reaching just level 0 (high-index $\mapsto$ level 0) is a repeater, has no devices, comes from a small size school with low parental involvement in the school running and bad school climate. (b) The student with the highest probability of reaching level 1 (low-index $\mapsto$ level 1) comes from better-off family (with positive SES index), has 4 books (2.5 books above mean), attend a private school with just 12% rate of disadvantaged students, and has 11 devices (3 above the mean country value for this sample).
  • Brazil. (a) Profile of student with achievement $\mapsto$ level 0: attend a school with full disadvantage, comes from low SES poor family, has zero books and, surprisingly, an educated father. (b) Profile of student with achievement $\mapsto$ level 1: does more regularly homework (above average), attend a school with below average index of bad school climate, satisfaction with life in the middle of the scale, a low educated mother and most likely to work for paid everyday.
  • Chile. (a) Profile of student with achievement $\mapsto$ level 0: is a repeater who also skipped school (1 whole day last 2 weeks), whose mother only completed lower secondary, attends a school of a size in the bottom quartile with good safety in its surroundings areas. (b) Profile of student with achievement $\mapsto$ level 1: attend a school located in a even safer area and with a even better school climate and with a relatively high contextual shortcomings, has only one sibling, and did not repeat primary.
  • Colombia. (a) Profile of student with achievement $\mapsto$ level 0: comes from a poor family (SES into the bottom half of distribution), has none digital devices and repeated primary, attend a school with quite large index for bad school climate, and whose mother is highly educated. (b) Profile of student with achievement $\mapsto$ level 1: comes from a better-off family and does not engage in paid-work, has 3 books (top 90%), and attend a school without disadvantaged students and with a high degree of autonomy (index large and positive).
  • Dominican Rep. (a) Profile of student with achievement $\mapsto$ level 0: repeated primary, attend a school with poor school climate, comes from low-SES family, is satisfied with life, and in the school there is a support of inclusivity. (b) Profile of student with achievement $\mapsto$ level 1: comes also from a school with poor school climate but with the school located in safer area, the extent of support of inclusivity in the school is larger, does not work for pay and is satisfied with life.
@@
For the second set of five countries (Table 7), main findings are:
  • Mexico. (a) Profile of student with achievement $\mapsto$ level 0: is engaged fully in paid work, does not have any books, missed school (3 months in a row), and repeated primary. (b) Profile of student with achievement $\mapsto$ level 1: comes from relatively poor family with low SES index, but is not engaged in paid work and does regularly homework, attend a school located in an unsafe area, and in a school with parents have low participation.
  • Panama. (a) Profile of student with achievement $\mapsto$ level 0: attend a highly disadvantaged school with a lack of PCs connected to the internet, does hardly any homework, comes from really poor household whose mother did not complete ISCED level 1 (primary). (b) Profile of student with achievement $\mapsto$ level 1: comes from a family whose SES is in the top quartile and has 9 digital devices (large than the median by 2), though the school he//she attends has very low autonomy and no qualified teachers.
  • Peru. (a) Profile of student with achievement $\mapsto$ level 0: comes from an Indigenous background who repeated primary and does just less than 1 hour homework per week, and attends a school with severe low school climate and very weak ICT infrastructure. (b) Profile of student with achievement $\mapsto$ level 1: does not work for payment, has 8 digital devices at home (above median), the family wealth falls into the top half of the SES distribution, an attend school with a STR of 32 with high autonomy.
  • Paraguay. (a) Profile of student with achievement $\mapsto$ level 0: comes from a household with 3 devices and high SES, only engage 1 day in unpaid work, but attend a school with low autonomy and just 16% of teacher attending PD. (b) Profile of student with achievement $\mapsto$ level 1: is richer than a level 0 student (higher value of SES index) and has more devices at home, attend a private school with larger autonomy of a total size of $\approx$ 900 students.
  • Uruguay. (a) Profile of student with achievement $\mapsto$ level 0: primary repeater, low life satisfaction, no devices at home, whereas (b) Profile of student with achievement $\mapsto$ level 1: did not repeat primary, family SES is larger (in the 3rdrd quartile), and has 3 books at home (median for this data is just 1), attends a school with more than 70% of either certified teachers or teachers attending PD.
@@ 6 Discussion
Latin American students’ limited achievement at secondary level, underpinned by systemic embedded inequalities, are very concerning. This inequality-driven learning crisis has been made worse by the COVID-19 pandemic (Betthäuser et al., 2023; UNESCO, 2024), with the region experiencing the larger period of school closures (of around 269 days, Bracco et al., 2025; Gómez and Andrés Uzín P, 2022). The region learning crisis is severe, with only between 5 and 7 out of 10 students reaching adequate competency in reading and math (Arias Ortiz et al., 2024b). Is within this bleak scenario that the paper’s contribution should be understood, namely, the profiling of low achievers, so that adequate interventions can ameliorate the extent of the learning crisis in the region.
@@
In particular, I follow a mechanism to identify poor performing students through explainable ML methods (Chen et al., 2022; Lundberg and Lee, 2017; Štrumbelj and Kononenko, 2014), i.e., Shapley values. The paper contributes to the emerging body of studies using interpretable ML results in education (Ali et al., 2025; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022; Zhu et al., 2025), though chiefly, expanding this literature to global south education systems at a regional scale.
@@
The paper, relying equally in global and local explanations at a regional level and per country, yields several insights into the LAC region learning crisis by offering a detailed account of which factors may sturdily place a students in the lowest bracket of performance (level 0, no competency achieved) or, alternative, which malleable factors may allow a bottom performer to transition to the next achievement bracket, level 1. The same analysis is carried out for the level 1 to level 2 students (where in level 2 basic subjects competencies are achieved). This granular analysis at the bottom of the learning distribution is fundamental for successful targeted interventions pushing up students to the next level of achievement. In essence, the paper’s approach and its policy implications are close to the existing literature of the region for the detection of secondary students at risk of dropout (e.g., Haimovich and Adelman, 2021; Rodríguez et al., 2023; Vinas-Forcade et al., 2021) via the setup of automatic alert systems (Early Warming Systems).
@@
The utility of Local SHAP values cannot be undervalued. They are a powerful policy tool as they offer an identification strategy of students locked into lower performing groups (or at risk of falling into that group) because they, not only rank leading covariates revealing positive or negative impacts of each variable on the model output for each observation (student), but also because they give the specific values each covariates take in that ranking, thereby given precise profiles of students. Results from the SHAP method can further guide follow-up qualitative studies attempting to gather more in-depth information on the reasons behind poor performance, a sequential combination of ML approaches and interviews used in late research (Gonzalez, 2021; Jayachandran et al., 2023). The sequential application of estimated SHAP values in learning surveys with qualitative interviews (mixed methods), is an interesting avenue for future research which can offer more insights into the learning crisis problem. Below, I flesh out policy implications from leading regional estimates which can be framed as a gradual two-way process of interventions: (i) first, to identify those at the bottom of achievement; and (ii) to introduce leverage mechanisms to modify covariates of level 0 (or level 1) students towards profiles of level 1 (level 2) students.
@@
In the local regional analysis, I find that the profile for a student who is locked-into the lack of competency group (level 0) is Indigenous, a repeater, is engaged in paid work (half of the week) and comes from a very poor household without digital devices. Thus, an entry point of a policy is to target poor Indigenous students and//or poor repeaters with financial support so that to avoid (or minimise) the impact on learning of paid work, likely leading to repetition. Culture responsive programs on bilingual education, raising Indigenous agency (López, 2021) and the attraction of good teachers to schools with mostly Indigenous students (Cortina, 2017; Delprato et al., 2022), tackling also repetition earlier on in primary (see: Learning Spaces program in Dominican Rep; Chávez et al., 2021), and conditional cash transfers (e.g., see examples of Brazil and Peru Draeger, 2021; Gaentzsch, 2020) alleviating disadvantaged students financial constraints, would all be beneficial interventions for level 0 students.
@@
Conversely, I find that students performing at level 1 do not work, are not repeaters, have a higher number of books and digital devices at home. Hence, shifting leading covariates from the profile of level 0 performing student to a level 1 performing student, would also require additional public investment on educational inputs. This shift, in turn, is related to programs such as the one laptop books in the region (e.g in rural schools in Peru: Cristia et al., 2017, or in Uruguay: Díaz et al., 2022; Melo et al., 2017). Note as well that, improving school climate and boosting students’ well-being support from teachers, are features of level 1 student profile which can benefit the lowest level 0 performers. The overall importance of school climate for learning is in line with the recent reviews of Varela et al. (2024) and Chávez et al. (2021), and the need of targeting boys of public schools in the region as an initial step (Delprato et al., 2017).
@@
Besides, a comparison of school features between these two level 0 and level 1 students’ profiles —and as found by earlier studies (Agasisti et al., 2023; Delprato, 2025; Mereles and Canese, 2022)— highlights the need of higher investments in schools’ ICT stock and connectivity, and promoting the collaboration on the use of digital devices by teachers (Arias Ortiz et al., 2024a). As shown by earlier literature, weak poor ICT access in the LAC region is indeed a barrier for those students attempting to move up in the learning distribution towards performance levels where subjects competencies are in fact reached. For instance, CEPAL (2020)’s report indicates that only 70% of people in the LAC region poses internet connection, and just 46% of children (aged 5-12) live in households without an internet connection. High gaps on ICT access by socioeconomic status are also observed, with students in the top wealth quartile being at least four times as likely (in comparison to the students from the bottom quartile) to have a laptop at home (Berlanga et al., 2020). This comparison also reveals that more investment is required on teachers preparation and teachers’ numbers in pedagogic roles (Ainscow et al., 2025; Miranda et al., 2025), ensuring quality and promoting continuous professionalisation.
@@
Similarly, policies seeking out to move up students from level 1 to level 2 would need to focus on —as well as enhancing level 1’s students stock of educational inputs and alleviating poverty and the chances of being involved in paid work—, offering better conditions for level 1 students to be able to carry out more homework (e.g., after school classes), and boosting family educational support. At the school level, policies should reshape schools’ functioning on different dimensions: boost involvement on parents on schools’ governance, and to enhance teaching quality too (via attracting teachers having more certification and attendance of PD training).
@@
The analysis has some limitations. Methodologically, a more exhaustive approach for benchmarking students profiles would be helpful, producing a more robust contrasting of students’ profiles, based on incremental steps (e.g., deciles) of the SHAP values index $\big[\Phi^{\text{min}}_{i}$, $\Phi^{\text{q10}}_{i}$$\Phi^{\text{q90}}_{i}$, $\Phi^{\text{max}}_{i}\big]$. Rather than focusing on two students for each sample, a continuum of profiles along students’ distribution of SHAP values would allow a better comparison on shift of covariates underpinned by future scaffolding interventions. Also, the use of SHAP values has found some methodological critics (Kumar et al., 2020), though this critic is an overall problem for model explanation methods (Chen et al., 2023). A future direction of research is to address intersectionality (combinations of student//family covariates with school covariates) which is a driving concept behind how inequality appears and persist in resource-constrained education systems (Codiroli Mcmaster and Cook, 2019). I succinctly address this in Appendix E (Figure E1) with SHAP values interactions, though further work in this direction is needed so that interventions are designed using joint markers of personal and contextual disadvantages from students and schools.

### v2

@@ 5.4 Countries’ results
Here we examine SHAP values at the country level by running a similar analysis per country for the level 0 and level 1 sample13 yielding local explanations of SHAP values. But, first, we present the global contributions, i.e., the average SHAP (absolute) values for all covariates.14 As a summary, we plot the frequency of the highest SHAP values contributions on predictions across countries. In Figure 6 we select the top 10 ranked covariates for each country and then obtain the number of times a given covariate appears in the top 10 list. The figure shows that primary repetition appears in the top ranked list for all 10 countries, while family SES and school climate (in 9 countries) and digital devices (in 8 countries). This is followed by lack of school safety and homework intensity (in 6 countries) whereas the covariates paid work, mother education and number of books appears in the top 10 ranked covariates for 5 countries.
@@
As regards to local explanations, the main takeaways concerning leading drivers for students more likely to either reach level 0 or level 1 achievement per country are outlined below.15 For the first set of five countries (Table 6):
  • Argentina. (a) The student with the largest probability of reaching just level 0 (high-index $\mapsto$ level 0) is a repeater, has no devices, comes from a small size school with low parental involvement in the school running and bad school climate. (b) The student with the highest probability of reaching level 1 (low-index $\mapsto$ level 1) comes from a better-off family (with positive SES index), has 4 books (2.5 books above mean), attends a private school with just 12% rate of disadvantaged students, and has 11 devices (3 above the country’s mean).
  • Brazil. (a) Profile of student with achievement $\mapsto$ level 0: attends a school with full disadvantage, comes from low SES family, has zero books. (b) Profile of student with achievement $\mapsto$ level 1: does more regularly homework (above average), attends a school with below average index of bad school climate, has a satisfaction with life in the middle of the scale, has a low educated mother and is likely to be engaged in paid work everyday.
  • Chile. (a) Profile of student with achievement $\mapsto$ level 0: is a repeater who also skipped school (1 whole day last 2 weeks), whose mother only completed lower secondary, attends a school of a size in the bottom quartile with good safety in its surroundings areas. (b) Profile of student with achievement $\mapsto$ level 1: attends a school located in a even safer area and better school climate, has one sibling, and did not repeat primary.
  • Colombia. (a) Profile of student with achievement $\mapsto$ level 0: comes from a poor family (SES into the bottom half of distribution), has none digital devices and repeated primary, attends a school with bad school climate, and has a mother who is educated. (b) Profile of student with achievement $\mapsto$ level 1: comes from a better-off family and does not engage in paid-work, has 3 books (top 90%), and attends a school without disadvantaged students and with a high degree of autonomy (index large and positive).
  • Dominican Rep. (a) Profile of student with achievement $\mapsto$ level 0: repeated primary, attends a school with poor school climate, comes from low-SES family, is satisfied with life, and in the school there is a support of inclusivity. (b) Profile of student with achievement $\mapsto$ level 1: comes also from a school with poor school climate but with the school located in safer area, the extent of support of inclusivity in the school is larger, does not work and is satisfied with life.
@@
For the second set of five countries (Table 7), main findings are as follows:
  • Mexico. (a) Profile of student with achievement $\mapsto$ level 0: is engaged fully in paid work, does not have any books, missed school (3 months in a row), and repeated primary. (b) Profile of student with achievement $\mapsto$ level 1: comes from relatively poor family with low SES index, but is not engaged in paid work and does regularly homework, attends a school located in an unsafe area, and in a school where parents have low participation.
  • Panama. (a) Profile of student with achievement $\mapsto$ level 0: attends a highly disadvantaged school with a lack of PCs connected to the internet, does hardly any homework, comes from really poor household whose mother did not complete primary. (b) Profile of student with achievement $\mapsto$ level 1: comes from a family whose SES is in the top quartile and has 9 digital devices (large than the median by 2), though the school he//she attends has very low autonomy and no qualified teachers.
  • Peru. (a) Profile of student with achievement $\mapsto$ level 0: comes from an Indigenous background who repeated primary and does just less than 1 hour homework per week, and attends a school with severe low school climate and very weak ICT infrastructure. (b) Profile of student with achievement $\mapsto$ level 1: does not work for payment, has 8 digital devices at home (above median), the family wealth falls into the top half of the SES distribution, an attend school with a STR of 32 and high autonomy.
  • Paraguay. (a) Profile of student with achievement $\mapsto$ level 0: comes from a family with 3 devices and high SES, only engages 1 day in unpaid work, but attend a school with low autonomy which has just a rate of teacher attending PD of 16%. (b) Profile of student with achievement $\mapsto$ level 1: is richer than a level 0 student (higher value of SES index) and has more digital devices at home, attends a private school with larger autonomy, and a school’s enrolment of $\approx$ 900 students.
  • Uruguay. (a) Profile of student with achievement $\mapsto$ level 0: is a primary repeater, has low life satisfaction and no devices at home. (b) Profile of student with achievement $\mapsto$ level 1: did not repeat primary, family SES is larger (in the 3rd quartile), has 3 books at home (median for this covariate is just 1), attends a school with more than 70% of either certified teachers or teachers attending PD.
@@
5.5 Heterogeneity’s results
In this section we briefly address some aspects of intersectionality of the whole sample estimates. Namely, we show how estimations for the relationship of two key drivers behind low and bottom achievement of students (that is, family SES and parental educational input) vary by students’ gender, repetition and ICT stock.16 Interaction effects are derived using the term $\sum_{j\neq i}\Phi_{i,j}(f,x)$ from Eq. (9). Results are displayed in Figure 7 (for the bottom performers comparison) and in Figure 8 (for the low performers comparison).
On the one hand, Figures 7(a), 7(b) and 7(c) describe the interaction of family SES with a given student’s feature: gender, primary repetition and with ICT stock at home (i.e., number of digital devices). When the effect of family SES is compounded with gender (with blue dots denoting female students and red dots male students) it can be seen decreasing SHAP values along the wealth index for male students and increasing for female students; in other words, coming from a poorer household affects relative more a male student than female student and, conversely if a student stems from a better-off household, and without important gender gaps for students in the middle of the wealth index (Figure 7(a)). Specifically, the plot shows gaps on SHAP values of around +4 (and so a higher chances to fall into level 0 achievement group) for a female student when family SES is equal to 1 (a richer family) and, conversely, gaps on SHAP values of +6 for quite poor families (wealth index = -3) with male students being much more likely to be at the bottom of the learning distribution. In the case of primary repetition (Figure 7(b)) there is clear pattern about how wealth compounds the effect of repetition: poorer and repeater students have positive SHAP values (higher chances to be level 0 performer), with gaps on SHAP values against a poor but not repeater of around +15. Figure 7(c) shows that, when family SES is interacted with a student’s ICT stock, a poorer student with low number of digital devices (in blue) has slightly larger SHAP values whilst a richer student with larger number of devices (in red at wealth = + 1) has negative SHAP values of -0.1, with better chances to fall into the better performing group (level 1). On the other hand, estimated SHAP values for the interactions with family educational support have not much of a differential effect in the case of ICT stock (Figure 7(e)), though they do for gender. Figure 7(d) shows that lack of family educational support affects negatively female students (as most blue dots are above zero at values of the family support index << 0), but they don’t for male students.
[Figure 7 here]
Figure 8 contains estimates for SHAP values interactions for the level1-level 2 sample comparison. The gender gap is more marked here when the comparison moves up in the learning distribution. Household poverty impacts relatively more male students, particular for very poor families (index below -2) as gaps with female students SHAP values is $\approx$ +0.10, although from quite a long range and family disadvantage (wealth index $\geq$ -2) all female students are more prone to fall into level 1 (rather than the reaching proficiency at level 2) (Figure 8(a)). Primary repetition interaction with wealth (Figure 8(b)) does not seem to significant impact for a student falling into level 1 as SHAP interactions (red dots) hovers around zero. Along the same vein, interactions of ICT with both family SES and family educational support (Figures 8(c) and 8(e), respectively) and their estimated SHAP values are not different by level of ICT (or number of digital devices), thereby suggesting a lack of intersectionality among these factors. Yet, for gender and family educational support (Figure 8(d)), a distinct pattern on SHAP interactions values emerges boosting gender gaps (in favour of female students) at higher degree of family support, and in the reverse direction (detrimental for female students) at lower educational family support levels.
[Figure 8 here]
@@ 6 Discussion
Underpinned by broad inequalities in the education sector, Latin American students’ limited achievement at secondary level is concerning. This inequality-driven learning crisis has been made worse by the COVID-19 pandemic (Betthäuser et al., 2023; UNESCO, 2024), with the region experiencing the larger period of school closures (Bracco et al., 2025; Gómez and Andrés Uzín P, 2022). The region learning crisis is severe, with only between 5 and 7 out of 10 students reaching adequate competency in reading and math (Arias Ortiz et al., 2024b). Is within this bleak scenario that the paper’s contribution should be understood, namely, through the profiling of low achievers, so that adequate interventions can be put forward to mitigate the extent of the learning crisis in the region.
@@
In particular, this study follows a mechanism to identify poor performing students through explainable ML methods (Chen et al., 2022; Lundberg and Lee, 2017; Štrumbelj and Kononenko, 2014), i.e., Shapley values. The paper contributes to the emerging body of studies using interpretable ML results in education (Ali et al., 2025; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022; Zhu et al., 2025) and broadly to the literature about the identification of at-risk students (Shi et al., 2025, 2024); though, expanding this literature to global south education systems at a regional scale.
@@
The paper, equally relying on global and local explanations at a regional level and at a country level, yields several insights into the LAC region learning crisis by offering a detailed account of which factors may sturdily place a students in the lowest bracket of performance (level 0, no competency achieved) or, alternative, which malleable factors may allow a bottom performer to transition to the next achievement bracket (level 1). The same analysis is carried out for the level 1 and level 2 students (where level 2 performance denotes students achieving basic subjects competencies). This granular analysis at the bottom of the learning distribution is fundamental for successful targeted interventions pushing up students to the next level of achievement. In essence, the paper’s approach and its policy implications are close to the existing literature of the region for the detection of secondary students at risk of dropout (e.g., Haimovich et al., 2021; Rodríguez et al., 2023; Vinas-Forcade et al., 2021) via the setup of automatic alert systems (Early Warming Systems).
@@
The utility of local SHAP values cannot be underestimated. They are a powerful policy tool as they offer an identification strategy of students locked into lower performing groups (or at risk of falling into that group) because they, not only rank leading covariates revealing positive or negative impacts of each variable on the model output for each observation (student), but also because they give the specific values each covariates take in that ranking, thereby given precise profiles of students. Results from the SHAP method can also further guide follow-up qualitative studies attempting to gather more in-depth information on the reasons behind poor performance. This sequential combination of ML approaches and interviews has been applied in certain studies (e.g., Gonzalez, 2021; Jayachandran et al., 2023). The sequential application of estimated SHAP values in learning surveys with qualitative interviews is an interesting avenue for future research by offering more insights into the learning crisis problem. Below, we flesh out some policy implications from leading regional estimates which can be framed as a gradual two-way process of interventions: (i) first, to identify those at the bottom of achievement; and (ii) to introduce leverage mechanisms to modify covariates of level 0 (level 1) students towards profiles of level 1 (level 2) students.17
@@
In the local regional analysis we find that the profile for a student who is locked-into the lack of competency group (level 0) is Indigenous, a repeater, is engaged in paid work (half of the week) and comes from a very poor household without digital devices. Thus, an entry point of a policy is to target poor Indigenous students and//or poor repeaters with financial support so that to minimise the impact on learning of paid work, likely leading to repetition. Culture responsive programs on bilingual education, raising Indigenous agency (López, 2021) and the attraction of good teachers to schools with mostly Indigenous students (Cortina, 2017; Delprato et al., 2022), tackling repetition earlier on in primary (see: Learning Spaces program in Dominican Rep; Chávez et al., 2021), and cash transfers (see examples for Brazil and Peru Draeger, 2021; Gaentzsch, 2020) alleviating disadvantaged students financial constraints, would all be beneficial interventions for level 0 achieving students.
@@
Conversely, we find that students performing at level 1 do not work, are not repeaters, have a higher number of books and digital devices at home. Hence, shifting leading covariates from the profile of level 0 performing student to a level 1 performing student would also require additional public investment on educational inputs. This shift, in turn, is related to programs such as the one laptop and investment in books in the region (e.g in rural schools in Peru: Cristia et al., 2017, or in Uruguay: Díaz et al., 2022; Melo et al., 2017). Also, improving school climate and boosting students’ well-being support from teachers, are features of level 1 student profile which can benefit a lowest level 0 performer. The overall importance of school climate for learning is in line with the reviews of Varela et al. (2024) and Chávez et al. (2021).
@@
Besides, a comparison of school features between the two level 0 and level 1 students’ profiles–and as found by earlier studies (Agasisti et al., 2023; Delprato, 2025; Mereles and Canese, 2022)–highlights the need of higher investments in schools’ ICT stock and connectivity, and promoting the collaboration on the use of digital devices by teachers (Arias Ortiz et al., 2024a). Poor regional ICT access is a sturdy barrier for those students attempting to move up in the learning distribution towards performance levels where subjects competencies are in fact reached. For instance, CEPAL, 2020’s report indicates that only 70% of people in the LAC region poses internet connection and 46% of children (aged 5-12) live in households without an internet connection. High gaps on ICT access by socioeconomic status are also observed, with students in the top wealth quartile being at least four times as likely (in comparison to the students from the bottom quartile) to have a laptop at home (Berlanga et al., 2020). This comparison also reveals that more investment is required on teachers preparation and teachers’ numbers in pedagogic roles (Ainscow et al., 2025; Miranda et al., 2025), ensuring quality and promoting continuous professionalisation.
@@
Policies seeking out to move up students from level 1 to level 2 would need to focus on (as well as on enhancing level 1’s students stock of educational inputs and alleviating poverty and so the chances of being involved in paid work) offering better conditions for level 1 students to be able to carry out more homework (e.g., after school classes), and boosting family educational support. At the school level, policies should reshape schools’ functioning on different dimensions: develop involvement of parents on schools’ governance and to enhance teaching quality too (via attracting teachers who have more certification and higher attendance of PD training).
@@
The analysis has some limitations. Methodologically, a more exhaustive approach for benchmarking students profiles would be helpful, producing a more robust contrasting of students’ profiles, based on incremental steps (e.g., deciles) of the SHAP values index $\big[\Phi^{\text{min}}_{i}$, $\Phi^{\text{q10}}_{i}$$\Phi^{\text{q90}}_{i}$, $\Phi^{\text{max}}_{i}\big]$. Rather than focusing on two students for each sample, a continuum of profiles along students’ distribution of SHAP values would allow a better comparison on shift of covariates underpinned by future scaffolding interventions. A future direction of research is to address intersectionality (combinations of student//family covariates with school covariates) which is a driving concept behind how inequality appears and persist in resource-constrained education systems (Codiroli Mcmaster and Cook, 2019; Delprato, 2024). We succinctly address this with SHAP values interactions, though further work in this direction is indispensable so that interventions are designed using joint markers of personal and contextual disadvantages from students and schools.